A 40 foot beam is to be cut in two, if one portion is one third the length of the other portion what are the two lengths of the portions?
(1/3)(40/1)
Multiply (1*40) / (3*1)
=40/3
But now I'm stuck, help please? Have to use 5 steps process

Answer by sudhanshu_kmr(1152)


Assume length of the portions are x and 3x.
x + 3x = 40
=> 4x= 40
=> x = 40/4 = 10
so, length of the portions are 10 foot and 30 foot.